# Diffusion Non-Additive Model for Multi-Fidelity Simulations with Tunable Precision

### v1

@@ 1 Introduction
Figure 1: The visualizations of FEM solutions of the Poisson’s equation at $x=0.5$ for mesh sizes $t=0.09$, $0.07$, $0.05$, and $0.03$.
@@
A wide range of multi-fidelity modeling approaches has been developed in the literature. Notable examples include Kennedy and O’Hagan, (2000), Qian et al., (2006), Qian and Wu, (2008), Le Gratiet and Garnier, (2014), Perdikaris et al., (2017), Kerleguer et al., (2024), and Heo and Sung, (2025), along with work on experimental design (Yuchi et al.,, 2023; Sung et al.,, 2024) and optimization (Picheny et al.,, 2013; He et al.,, 2017). However, most of these methods focus on emulating the highest fidelity simulator available, often without explicitly modeling the tuning parameter that controls fidelity. Exceptions include the work by Tuo et al., (2014), which targets the exact solution in the limit of zero mesh size, and its adaptive kernel extension by Boutelet and Sung, (2025).
@@ 2.2 Diffusion Non-Additive (DNA) Model
where $t_{1}>t_{2}>\cdots>t_{L}>0$. This recursive formulation satisfies the Markov property, as each fidelity level $f_{l}(x)$ depends only on the immediately preceding level $f_{l-1}(x)$, without direct dependence on earlier levels.
The structure suggests a diffusion-like process, where fidelity levels evolve through a recursive relationship. The term “diffusion” is inspired by its analogy in deep learning models (Ho et al.,, 2020), where similar recursive structures propagate information across layers to enable progressive refinement. The hierarchical structure of the model is visualized in the top panel of Figure 2.
$f_{1}(\mathbf{x})$$\dots$$f_{l-1}(\mathbf{x})$$f_{l}(\mathbf{x})$$\dots$$f_{L}(\mathbf{x})$$f(t,\mathbf{x})$$W(t_{l},\mathbf{x},f_{l-1}(\mathbf{x}))$$W(t,\mathbf{x},f_{L}(\mathbf{x}))$$\mu^{*}_{1}(\mathbf{x})$$\dots$$\mu^{*}_{l-1}(\mathbf{x})$$\mu^{*}_{l}(\mathbf{x})$$\dots$$\mu^{*}_{L}(\mathbf{x})$$\mu^{*}(t,\mathbf{x})$$h_{1}(\mathbf{x}|t_{l},\mu^{*}_{l-1}(\mathbf{x}),\sigma^{*2}_{l-1}(\mathbf{x}))$$h_{1}(\mathbf{x}|t,\mu^{*}_{L}(\mathbf{x}),\sigma^{*2}_{L}(\mathbf{x}))$$\sigma^{*2}_{1}(\mathbf{x})$$\dots$$\sigma^{*2}_{l-1}(\mathbf{x})$$\sigma^{*2}_{l}(\mathbf{x})$$\dots$$\sigma^{*2}_{L}(\mathbf{x})$$\sigma^{*2}(t,\mathbf{x})$$h_{2}(\mathbf{x}|t_{l},\mu^{*}_{l-1}(\mathbf{x}),\sigma^{*2}_{l-1}(\mathbf{x}))$$h_{2}(\mathbf{x}|t,\mu^{*}_{L}(\mathbf{x}),\sigma^{*2}_{L}(\mathbf{x}))$ModelPosteriorMeanPosteriorVariance
@@
where $\delta_{l}(\mathbf{x})$ is unknown discrepancy function that converges to 00 as $t_{l}\to 0$.
@@
It is also worth noting that, unlike the recent Recursive Non‑Additive (RNA) model proposed by Heo and Sung, (2025), which constructs a separate GP at each fidelity level $l$ with its own hyperparameters, the DNA model incorporates the tuning parameter $t_{l}$ as an additional input and embeds all non-lowest‑fidelity outputs into a single GP. By augmenting the input space to $(t,\mathbf{x},y)$, DNA employs a shared lengthscale and scale parameter across levels $2,\dots,L$, reducing the total number of lengthscale hyperparameters from $(dL+L-1)$ in RNA to $(2d+1)$. Moreover, the explicit inclusion of $t$ facilitates DNA to perform direct inference for higher-fidelity outputs (i.e., extrapolation to smaller tuning parameters and even the exact solution corresponding to $t=0$).
@@ 2.3 Parameter Estimation
$-\frac{1}{2}\log(\det(\mathbf{K}))+\frac{N_{-1}}{2}\log\left((Y_{-1}-\hat{\alpha}\mathbf{1}_{N_{-1}})^{T}\mathbf{K}^{-1}(Y_{-1}-\hat{\alpha}\mathbf{1}_{N_{-1}})\right).$ (9)
@@ 3 Posterior of DNA Model
where $\mathbf{k}_{1}(\mathbf{x})$ is an $n_{1}\times 1$ matrix with elements $(\mathbf{k}_{1}(\mathbf{x}))_{i,1}=K_{1}(\mathbf{x},\mathbf{x}^{[1]}_{i})$.
@@
×$\times\left.\exp{\left(-\sum_{j=1}^{d}\frac{c_{i}(x_{j}-(\mathbf{X}_{-1})_{ij})^{2}+c_{k}(x_{j}-(\mathbf{X}_{-1})_{kj})^{2}}{\theta_{j}}\right)}\right),$
@@
×$\times\exp{\left(-\frac{c_{i}\left((Y_{-L})_{i}-\mu\right)^{2}+c_{k}\left((Y_{-L})_{k}-\mu\right)^{2}+\frac{2}{\theta_{y}}c_{i}c_{k}\sigma^{2}\left((Y_{-L})_{i}-(Y_{-L})_{k}\right)^{2}}{\theta_{y}+2(c_{i}+c_{k})\sigma^{2}}\right)}.$
@@
Proposition 3.1 enables efficient computation of the posterior mean and variance through a recursive formulation. The hierarchical structure of the posterior inference process is illustrated in Figure 2. To approximate the posterior distribution, we adopt the moment-matching method, leveraging the Gaussian approximation based on the derived posterior mean and variance. The unknown parameters in the posterior distribution are replaced by their corresponding estimates. An R package, DNAmf, implementing the proposed methods is available on an open repository.
@@ 4 Numerical Studies
where $r_{w}\in[0.05,0.15]$, $r\in[100,50000]$, $T_{u}\in[63070,115600]$, $H_{u}\in[990,1110]$, $T_{l}\in[63.1,116]$, $H_{l}\in[700,820]$, $L\in[1120,1680]$, and $K_{w}\in[9855,12045]$. The first one-dimensional function is from Tuo et al., (2014), which exhibits an additive structure (referred to as the additive function) and yields the exact solution $f(0,x)=\exp{(-1.4x)}\cos{(3.5\pi x)}$. The second function is adapted from Higdon, (2002) (referred to as the non-additive function), where $t$ is in the denominator, showing a nonlinear relationship between different tuning parameters. The third synthetic example is the multi-fidelity Currin function (Currin,, 1988) from Sung et al., (2024). The last synthetic example is the multi-fidelity borehole function adapted from Xiong et al., (2013). The input locations are generated from the nested space-filling design introduced by Le Gratiet and Garnier, (2014), at equally spaced tuning parameter $t$ over the ranges given in the last column of Table 1, which also summarizes the sample sizes and input dimensions for each function.
$n_{1}$ $n_{2}$ $n_{3}$ $n_{4}$ $n_{5}$ $n_{6}$ $d$ $t$
Additive 13 10 7 4 1 1 [1,3]
Non-additive 13 10 7 4 1 1 [0.5,2.5]
Currin 19 16 13 10 7 4 2 [0.1,0.6]
Borehole 35 30 25 20 15 10 8 [0.5,3]
Table 1: Sample sizes $n_{l}$, input dimension $d$, and the range of the tuning parameter $t$ for each synthetic example.
We compare the predictive performance of the proposed model (labeled DNAmf) with four existing methods: two that aim to emulate highest fidelity output $f(t_{L},\mathbf{x})$, including the Recursive Non-Additive emulator (RNAmf) by Heo and Sung, (2025) and the auto-regressive model (CoKriging) by Le Gratiet and Garnier, (2014), and other two that target the exact solution $f(0,\mathbf{x})$, including the nonstationary GP model by Tuo et al., (2014) using Brownian motion kernel function (BM), and its extension using fractional Brownian motion kernel function (FBM) proposed by Boutelet and Sung, (2025). Due to structural limitations, the RNAmf and CoKriging models are restricted to using only three fidelity levels ($l=2,3,4$ for the additive and non-additive functions, and $l=4,5,6$ for the Currin and the Borehole function).
@@
Figure 3 illustrates model predictions along with confidence intervals for the additive and non-additive functions across different values of the tuning parameter $t$. By assuming that each fidelity level depends only on the immediately preceding level, the model effectively captures the evolving relationship as $t$ decreases to 00. Notably, the model accurately predicts the function at $t=0$ despite the absence of design points at this value. The confidence intervals mostly contain the true function, demonstrating the effectiveness of the uncertainty quantification. This extrapolation capability demonstrates the model’s practical utility in situations where data at very small values of $t$ are unavailable, which reduces the need for costly data collection while maintaining predictive accuracy.
@@
Notably, both RNAmf and CoKriging show relatively suboptimal performance. This is because they fit separate GPs at each fidelity level, relying only on the limited design points available at that level. As a result, they cannot utilize the highest fidelity level in the additive and non-additive functions, where only a single output is available. Furthermore, these methods focus solely on emulating the highest fidelity output $f(t_{L},\mathbf{x})$, without extrapolating to $f(0,\mathbf{x})$. In contrast, DNAmf, BM, and FBM leverage all available data across fidelity levels to build a unified emulator capable of extrapolation. This ability to pool information leads to better predictive performance, particularly when each level has a limited number of design points.
The proposed model also offers interpretability through the parameter $\beta$, which captures the strength of interaction between $\mathbf{x}$ and $t$. For the non-additive function, a stronger coupling is observed, reflected in a higher average estimated interaction parameter $\hat{\beta}=0.053$, compared to $\hat{\beta}=0.002$ for the additive function, $\hat{\beta}=0.004$ for the Currin function, and $\hat{\beta}=0$ for the borehole function.
@@ 5 Real Applications
In this section, we present three case studies to evaluate the predictive performance of the proposed approach in scenarios where obtaining an exact solution within finite time or cost is infeasible or difficult. Simulations are conducted at five fidelity levels, corresponding to five different mesh resolutions. As in Section 4, the designs are generated using the nested space-filling design of Le Gratiet and Garnier, (2014).
@@
  • Vibration of Square Plate: The second case study focuses on the fourth natural frequencies in hertz (Hz) of a square elastic plate of size $10\times 10\times 1$ (Li et al.,, 2020). Three material properties serve as input variables: Young’s modulus ($x_{1}\in[1\times 10^{11},4\times 10^{11}]$), Poisson’s ratio ($x_{2}\in[0.2,0.4]$), and mass density ($x_{3}\in[6\times 10^{3},9\times 10^{3}]$). Fidelity is again controlled by the finite-element mesh size $t$ (see Figure 5), which is selected at five equally spaced values in $[0.35,0.55]$. Because an analytical solution for the plate’s frequencies at infinite resolution is unavailable and high-resolution simulations are computationally expensive, we treat $t=0.3$ as the finest mesh for evaluating model performance. Simulations at this resolution are conducted at 100 uniform test input locations, with each run taking up to 8 minutes. The frequencies are computed via FEM using MATLAB.
@@
    where $u(x,s)$ denotes the temperature at spatial coordinate $x\in[0,L]$ and time $s\in[0,S]$, and $D$ denotes the thermal diffusivity coefficient. As $s$ approaches the target time $S$, the solution generally exhibits increasing complexity both in its spatial gradients and in the coupling between space and time. Consequently, direct integration from $s=0$ to $s=S$ on a fine time grid can become prohibitive. To address this, we introduce the tuning parameter $t=S-s$ that measures the temporal distance from the simulation time $s$ to the target time $S$. Early time-step solutions, (i.e., larger $t$, smaller $s$) provide inexpensive, low-fidelity information, while those obtained closer to the target time (i.e., smaller $t$, larger $s$) yield high-fidelity results at greater cost. To evaluate model performance, we compute the exact solution at $t=0$ at 100 uniform test input locations.
As in Section 4, model performance is evaluated using the RMSE and CRPS on 100 test points with 100 repetitions. The sample sizes, input dimensions, and the ranges of the tuning parameters for each case study, are summarized in Table 2.
$n_{1}$ $n_{2}$ $n_{3}$ $n_{4}$ $n_{5}$ $d$ $t$
Poisson 15 12 9 6 3 1 [0.01,0.09]
Plate 25 20 15 10 5 3 [0.35,0.55]
Heat 15 12 9 6 3 1 [0.1,0.5]
Table 2: Sample sizes $n_{l}$, input dimension $d$, and the range of the tuning parameter $t$ for each case study.
@@
As shown in Figure 6, DNAmf generally outperforms both FBM and BM across all case studies, with the exception of RMSE in the Poisson equation, where it is slightly less accurate. However, DNAmf achieves better CRPS in that case, indicating improved uncertainty quantification. The estimated interaction parameter $\hat{\beta}$ further highlights differences in coupling behavior. For both the Poisson and Plate problems, $\hat{\beta}\approx 0$, suggesting that the kernel becomes effectively separable, as discussed in Section 2. This implies that the input effect does not vary substantially across different values of the tuning parameter $t$. In contrast, for the Heat equation, we observe $\hat{\beta}\approx 1$, indicating strong interaction between inputs and the fidelity level. These patterns are also reflected visually in Figure S1.

### v2

@@ 1 Introduction
Figure 1: The visualizations of FEM solutions of the Poisson’s equation at $x=0.5$ for mesh sizes $t=0.03$, $0.05$, $0.07$, and $0.09$.
@@
A wide range of multi-fidelity modeling approaches has been developed in the literature. Notable examples include Kennedy and O’Hagan, (2000), Qian et al., (2006), Qian and Wu, (2008), Le Gratiet and Garnier, (2014), Perdikaris et al., (2017), Kerleguer et al., (2024), and Heo and Sung, (2025), along with work on experimental design (Yuchi et al.,, 2023; Sung et al.,, 2024) and optimization (Picheny et al.,, 2013; He et al.,, 2017). However, most of these methods focus on emulating the highest fidelity simulator available, often without explicitly modeling the tuning parameter that controls fidelity. Exceptions include the work by Tuo et al., (2014), which targets the exact solution in the limit of zero mesh size, as well as its adaptive kernel extension by Boutelet and Sung, (2025) and multi-dimensional tuning parameter generalization by Ji et al., (2024). More recently, Oates et al., (2025) extended Richardson extrapolation within a probabilistic framework, focusing on extrapolating the tuning parameters.
@@ 2.2 Diffusion Non-Additive (DNA) Model
where $t_{1}>t_{2}>\cdots>t_{L}>0$. This recursive formulation satisfies the Markov property, as each fidelity level $f_{l}(\mathbf{x})$ depends only on the immediately preceding level $f_{l-1}(\mathbf{x})$, without direct dependence on earlier levels.
The structure suggests a diffusion-like process, where fidelity levels evolve through a recursive relationship. The term “diffusion” is inspired by its analogy in deep learning models (Ho et al.,, 2020), where similar recursive structures propagate information across layers to enable progressive refinement. Each fidelity level $f_{l}(\mathbf{x})$ represents the simulator output under a coarser or less accurate configuration such as a relaxed solver tolerance or fewer iterations. As the tuning parameter $t_{l}$ increases, these choices accumulate approximation error, gradually biasing the output away from the exact solution $f(0,\mathbf{x})$ as illustrated in Figure 1. This degradation during the data generation procedure via FEM can be viewed as a forward diffusion across fidelity levels. On the other hand, the recursion $f_{l}(\mathbf{x})=W(t_{l},\mathbf{x},f_{l-1}(\mathbf{x}))$ exploits this one-step Markov dependence to refine predictions toward $f(0,\mathbf{x})$, resembling a backward diffusion process. The hierarchical structure of the model is visualized in the top panel of Figure 2.
[…]
@@
where $\delta_{l}(\mathbf{x})$ is unknown discrepancy function that converges to 0 as $t_{l}\to 0$.
@@
It is also worth noting that, unlike the recent Recursive Non‑Additive (RNA) model (Heo and Sung,, 2025) and the nonlinear auto-regressive GP model (Perdikaris et al.,, 2017), which constructs a separate GP at each fidelity level $l$ with its own hyperparameters, the DNA model incorporates the tuning parameter $t_{l}$ as an additional input and embeds all non-lowest‑fidelity outputs into a single GP. By augmenting the input space to $(t,\mathbf{x},y)$, DNA employs a shared lengthscale and scale parameter across levels $2,\dots,L$, reducing the total number of lengthscale hyperparameters from (dL+L1)(dL+L-1) in RNA to $(2d+1)$. Moreover, the explicit inclusion of $t$ facilitates DNA to perform direct inference for higher-fidelity outputs (i.e., extrapolation to smaller tuning parameters and even the exact solution, i.e., $t=0$).
@@ 2.3 Parameter Estimation
$\frac{1}{2}\log(\det(\mathbf{K}))+\frac{N_{-1}}{2}\log\left((Y_{-1}-\hat{\alpha}\mathbf{1}_{N_{-1}})^{T}\mathbf{K}^{-1}(Y_{-1}-\hat{\alpha}\mathbf{1}_{N_{-1}})\right).$ (9)
@@ 3 Posterior of DNA Model
[…]
@@
$\times\left.\exp{\left(-\sum_{j=1}^{d}\frac{c_{i}(x_{j}-(\mathbf{X}_{-1})_{ij})^{2}+c_{k}(x_{j}-(\mathbf{X}_{-1})_{kj})^{2}}{\theta_{j}}\right)}\right),$
@@
$\times\exp{\left(-\frac{c_{i}\left((Y_{-L})_{i}-\mu\right)^{2}+c_{k}\left((Y_{-L})_{k}-\mu\right)^{2}+\frac{2}{\theta_{y}}c_{i}c_{k}\sigma^{2}\left((Y_{-L})_{i}-(Y_{-L})_{k}\right)^{2}}{\theta_{y}+2(c_{i}+c_{k})\sigma^{2}}\right)}.$
@@
Proposition 3.1 enables efficient computation of the posterior mean and variance through a recursive formulation. The hierarchical structure of the posterior inference process is illustrated in Figure 2. To approximate the posterior distribution, we adopt the moment-matching method, leveraging the Gaussian approximation based on the derived posterior mean and variance. This Gaussian approximation to the DNA model minimizes the Kullback–Leibler divergence between the true distribution and its Gaussian representation (Minka,, 2001). The unknown parameters in the posterior distribution are replaced by their corresponding estimates. An R package, DNAmf, implementing the proposed methods is available on an open repository.
@@ 4 Numerical Studies
where $r_{w}\in[0.05,0.15]$, $r\in[100,50000]$, $T_{u}\in[63070,115600]$, $H_{u}\in[990,1110]$, $T_{l}\in[63.1,116]$, $H_{l}\in[700,820]$, $L\in[1120,1680]$, and $K_{w}\in[9855,12045]$. The first one-dimensional function is from Tuo et al., (2014), which exhibits an additive structure (referred to as the additive function) and yields the exact solution $f(0,x)=\exp{(-1.4x)}\cos{(3.5\pi x)}$. The second function is adapted from Higdon, (2002) (referred to as the non-additive function), where $t$ is in the denominator, showing a nonlinear relationship between different tuning parameters. The third synthetic example is the multi-fidelity Currin function (Currin,, 1988) from Sung et al., (2024). The last synthetic example is the multi-fidelity Borehole function adapted from Xiong et al., (2013).
The tuning parameter $t_{l}$ is chosen using a systematic refinement approach. Specifically, a geometric refinement is employed such that $t_{l}=ct_{l-1}$ with a constant ratio $c<1$, ensuring proportional reduction of the discretization spacing across levels. In scientific computing, such geometric refinement serves as a fundamental component in assessing the observed order of accuracy and the Grid Convergence Index, which formally verifies numerical convergence and quantifies discretization uncertainty (Oberkampf and Roy,, 2025). In the present numerical studies, we set $c=0.7$. The choice of sample sizes follows a similar geometric relation, given by $n_{l-1}=\lfloor c^{-\gamma}n_{l}\rceil$, where $\gamma=2$, which maintains a consistent balance between the refinement of the tuning parameter and the allocation of sample sizes across levels. Additional results under different choices of $c$ and $\gamma$ are provided in supplementary materials S7. The input locations are generated from the nested space-filling design introduced by Le Gratiet and Garnier, (2014), with the largest tuning parameter $t_{L}$ given in the last column of Table 1, which also summarizes the sample sizes and input dimensions for each function.
$n_{1}$ $n_{2}$ $n_{3}$ $n_{4}$ $n_{5}$ $n_{6}$ $d$ $t_{L}$
Additive 17 8 4 2 1 1 2.5
Non-additive 17 8 4 2 1 1 2.5
Currin 71 35 17 8 4 2 2 2.5
Borehole 71 35 17 8 4 2 8 2.5
Table 1: Sample sizes $n_{l}$, input dimension $d$, and the largest tuning parameter $t_{L}$ for each synthetic example.
We compare the predictive performance of the proposed model (labeled DNAmf) with four existing methods: two that aim to emulate highest fidelity output $f(t_{L},\mathbf{x})$, including the Recursive Non-Additive emulator (RNAmf) by Heo and Sung, (2025) and the auto-regressive model (CoKriging) by Le Gratiet and Garnier, (2014), and the nonlinear auto-regressive GP model (NARGP) by Perdikaris et al., (2017). The remaining two methods target the exact solution $f(0,\mathbf{x})$, namely the nonstationary GP model by Tuo et al., (2014) using Brownian motion kernel function (BM), and its extension using fractional Brownian motion kernel function (FBM) proposed by Boutelet and Sung, (2025). Due to structural limitations, the RNAmf, CoKriging, and NARGP models are restricted to using only three fidelity levels ($l=1,2,3$ for the additive, non-additive, and Borehole functions, and $l=3,4,5$ for the Currin function).
@@
Figure 3 illustrates model predictions along with confidence intervals for the additive and non-additive functions across different values of the tuning parameter $t$. By assuming that each fidelity level depends only on the immediately preceding level, the model effectively captures the evolving relationship as $t$ decreases to 0. Notably, the model accurately predicts the function at $t=0$ despite the absence of design points at this value. The confidence intervals mostly contain the true function, demonstrating the effectiveness of the uncertainty quantification. This extrapolation capability demonstrates the model’s practical utility in situations where data at very small values of $t$ are unavailable, which reduces the need for costly data collection while maintaining predictive accuracy.
@@
Notably, RNAmf, CoKriging, and NARGP show relatively suboptimal performance. This is because they fit separate GPs at each fidelity level, relying only on the limited design points available at that level. As a result, they cannot utilize the highest fidelity level in the additive and non-additive functions, where only a single output is available. Furthermore, these methods focus solely on emulating the highest fidelity output $f(t_{L},\mathbf{x})$, without extrapolating to $f(0,\mathbf{x})$. In contrast, DNAmf, BM, and FBM leverage all available data across fidelity levels to build a unified emulator capable of extrapolation. This ability to pool information leads to better predictive performance, particularly when each level has a limited number of design points.
The proposed model also offers interpretability through the parameter $\beta$, which captures the strength of interaction between $\mathbf{x}$ and $t$. For the non-additive function, a stronger coupling is observed, reflected in a higher average estimated interaction parameter $\hat{\beta}=0.007$, in contrast to $\hat{\beta}=0$ for the additive, Currin, and Borehole function.
@@ 5 Real Applications
In this section, we present three case studies to evaluate the predictive performance of the proposed approach in scenarios where obtaining an exact solution within a finite time or cost is infeasible or difficult. Simulations are conducted at five fidelity levels, corresponding to five different mesh resolutions. As in Section 4, the designs are generated using the nested space-filling design of Le Gratiet and Garnier, (2014).
@@
  • Vibration of Square Plate: The second case study focuses on the fourth natural frequencies in hertz (Hz) of a square elastic plate of size $10\times 10\times 1$ (Li et al.,, 2020). Three material properties serve as input variables: Young’s modulus ($x_{1}\in[1\times 10^{11},4\times 10^{11}]$), Poisson’s ratio ($x_{2}\in[0.2,0.4]$), and mass density ($x_{3}\in[6\times 10^{3},9\times 10^{3}]$). Fidelity is again controlled by the finite-element mesh size $t$ (see Figure 5). Because an analytical solution for the plate’s frequencies at infinite resolution is unavailable and high-resolution simulations are computationally expensive, we treat $t=0.3$ as the finest mesh for evaluating model performance. Simulations at this resolution are conducted at 100 uniform test input locations, with each run taking up to 8 minutes. The frequencies are computed via FEM using MATLAB.
@@
    where $u(x,s)$ denotes the temperature at spatial coordinate $x\in[0,L]$ and time $s\in[0,S]$, and $D$ denotes the thermal diffusivity coefficient. As $s$ approaches the target time $S$, the solution generally exhibits increasing complexity both in its spatial gradients and in the coupling between space and time. Consequently, direct integration from $s=0$ to $s=S$ on a fine time grid can become prohibitive. To address this, we introduce the tuning parameter t=Sst=S-s that measures the temporal distance from the simulation time $s$ to the target time $S$. Early time-step solutions, (i.e., larger $t$, smaller $s$) provide inexpensive, low-fidelity information, while those obtained closer to the target time (i.e., smaller $t$, larger $s$) yield high-fidelity results at greater cost. To evaluate model performance, we compute the exact solution at $t=0$ at 100 uniform test input locations.
As in Section 4, model performance is evaluated using the RMSE and CRPS on 100 test points with 100 repetitions. The sample sizes, input dimensions, the largest value of the tuning parameters, and constant ratio $c$ for each case study are summarized in Table 2. These mesh sizes and sample sizes are chosen in a similar fashion as in Section 4.
$n_{1}$ $n_{2}$ $n_{3}$ $n_{4}$ $n_{5}$ $d$ $t_{L}$ $c$
Poisson 17 11 7 5 3 1 0.1 0.65
Plate 28 21 15 11 8 3 0.55 0.9
Heat 25 15 9 5 3 1 0.5 0.7
Table 2: Sample sizes $n_{l}$, input dimension $d$, the largest tuning parameter $t_{L}$, and the constant $c$ that determines the other tuning parameters, $t_{l}=ct_{l-1}$, for each case study.
@@
As shown in Figure 6, DNAmf generally outperforms all other competitors across all case studies, with the exception of RMSE in the Poisson equation, where it is slightly less accurate than both FBM and BM. However, DNAmf achieves better CRPS in that case, indicating improved uncertainty quantification. Notably, RNAmf performs considerably worse for the Poisson equation, with both RMSE and CRPS values exceeding the displayed range in Figure 6. This degraded performance likely stems from its structural limitation of fitting separate GPs at each fidelity level using relatively few design points, which restricts its ability to effectively capture cross-level dependencies. The estimated interaction parameter $\hat{\beta}$ and the visualizations in Figure S5 together highlight differences in coupling behavior. For both the Poisson and Plate problems, $\hat{\beta}\approx 0$, suggesting that the kernel becomes effectively separable. This indicates that the influence of the input variables remains nearly constant across different values of the tuning parameter $t$, which is consistent with the top row of Figure S5, where model predictions and confidence intervals show little variation over $t$. In contrast, for the Heat equation, $\hat{\beta}\approx 1$, revealing the pronounced interaction between inputs and the fidelity level. This behavior is also evident in the bottom row of Figure S5, where the predictive results vary substantially with $t$, clearly reflecting the strong interaction between the input variables and tuning parameter captured by the nonseparable kernel.